A sample of nitrogen gas occupies a volume of 1.3 L at 836 mm Hg and 2° C. The volume increases by 1.1 L and the temperature decreases to 27 K. What is the final pressure exerted on the gas in mm Hg? Round to two decimal places.


1
Expert's answer
2021-02-28T06:15:28-0500

"V_1 = 1.3L\\
P_1 = 836mmHg
\\ T_1 = 2°C = 275K"


"V_2 = 2.4L
\\P_2 = ?

\\
T_2 =248K"


from,


"\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}"


"\dfrac{836×1.3}{275} = \dfrac{P_2×2.4}{248}"


"P_2 = \dfrac{836×1.3×248}{2.4×275} ="


408.37 mmHg.